OpenStudy (anonymous):

Help please x - y = 6 x = y + 2

OpenStudy (hexagon001):

solve simultaneously..

OpenStudy (hexagon001):

first rearrange the second so both the x and y are on one side

OpenStudy (anonymous):

(x=4, y=2}

OpenStudy (anonymous):

I TRIED THAT.

OpenStudy (anonymous):

X+y=2

OpenStudy (hexagon001):

rearranging the second equation entails taking the y to the left side... since it is positive..when we take to the other side it will become opposite to positive which is negative... so what do you think it is..

OpenStudy (anonymous):

or x-y=2

OpenStudy (hexagon001):

yes x-y=2

OpenStudy (hexagon001):

so now we have x-y=6 x-y=2 can you solve this simultaneously?

OpenStudy (anonymous):

0,6

OpenStudy (hexagon001):

nope.

OpenStudy (anonymous):

6,0

OpenStudy (hexagon001):

i take it you dont know how to solve simultaneuosly?

OpenStudy (anonymous):

No

OpenStudy (hexagon001):

ok lets take a different approach

OpenStudy (hexagon001):

in the second equation what does x equal to?

OpenStudy (anonymous):

2

OpenStudy (hexagon001):

no in the second equation x equals y+2... you must include everything after the equal sign...even if it be a letter do you get that?

OpenStudy (anonymous):

Yes

OpenStudy (hexagon001):

now substitute x=y+2 into the first equation

OpenStudy (hexagon001):

are you getting this?

OpenStudy (anonymous):

So IDK

OpenStudy (hexagon001):

in place of the x in the first equation put y+2

OpenStudy (anonymous):

Y+2=Y+2

OpenStudy (hexagon001):

no no no.. we have x-y=6.....first equation x=y+2.....second equation put y+2 in place of the x inthe first equation

OpenStudy (anonymous):

Y+2=x

OpenStudy (hexagon001):